OpenStudy (geekfromthefutur):

I will give medal and fan math isnt my best subject. Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (yanasidlinskiy):

I'm going to still help anyways:) P = the profit they make n = the number of flowers they sell m = the mark-up of the flowers c = the cost of a flower Then P = mcn - cn Rearrange to find n P = n(mc - c) n = P / (mc - c) Substituting values n = 500 / (1.4*0.5 - 0.5) n = 2500 They must sell 2500 flowers to make a profit of $500.

OpenStudy (yanasidlinskiy):

Do you understand it? Makes sense?

OpenStudy (geekfromthefutur):

Yeah its good thanks^-^

OpenStudy (yanasidlinskiy):

You're welcome!:) Glad I could help:) Even though you closed it, I still knew you needed help with it!:)

OpenStudy (geekfromthefutur):

:)Yup

OpenStudy (yanasidlinskiy):

lol..